40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1170     Entry time: Thu Nov 18 09:27:18 2010
Author: Koji 
Type: Laser 
Category: GYRO 
Subject: RFAM & Sprious RF coupling  
  • Checked the alignment of the beam at the input optics
  • Checked the RF leakage to the PD.
  • It turned out the isolation transformer and the cable at the function generator caused the main leakage source.
  • By removing them the leakage was improved 20~30dB.
  • At the end of the experiment the RFAM level (RF leakage + RFAM) was -85dBm for the beam of ~100mVDC
  • The alignment of the interferometer is now totally screwed up.

 1. Alignment check of the IO

The original angle reads of the wave plates were 316deg and 38.0deg for the QWP and HWP, respectively.
The beam power before the EOM was 284mW.

  • FIrst, the wave plates after the laser were adjusted so as to minimize the PBS transmission.
  • The PBS was also aligned to minimize the transmission.
  • With the angle setting of 318.2deg and 78.3deg, the power of 48uW reached the EOM.
     
  • Rotated the HWP to maximize the PBS transmission.

At the last state, with the angle of 318.2deg and 33.7deg, the power of 284mW reached the EOM.

  • Beam alignment to the EOM was checked. It was OK, but the slight touch of the alignment was made.

After the adjustment the power after the EOM was 274mW. The transmission efficiency of 96.5% sounds pretty fine.

2. RF leakage investigation

  • The primary cavity RF PD was aligned to the beam. I temporarily removed the attenuation mirror. The DC output (with 50ohm) was 90.9mV
  • I tried to minimize the 18MHz RFAM by tuning the HWP for the EOM.
  • I minimized the RFAM but I noticed that the amount of the RF is comparable with and without the beam
     
  • The leakage RF power at 18MHz was ~70dBm. This corresponds to the RIN of 10-3. It was there even without the beam on the PD.
     
  • I tracked down the cause of the leakage and found that the leakage power was reduced when the cable between the isolation transformer and the coupler is replaced to the SMA cable.
  • Another discovery was that the leakage decreased when the isolation transformer is removed.
  • See the attached photo 1.

  • The overall improvement was 20-30dB (seen in the attachment 2) although the peak height is continuouwly changing.
     
  • It is always difficult to understand how the RF leakage is improved: The isolation transformer isolate the grounding of the RF source and the others. The line is somehow grounded at various places. Maybe at the EOM or/and the LO ports? They may have not low impedance to the reference ground of our system and cause fluctuation of the voltage level on the table, the aluminum frame of the table, or the power line. Then they couples to the PD???
  • I imagine that we previously cancelled the RF leakage by putting an artificial RFAM by the EOM. But both the amount of the leakage and the RFAM was always changing. They may have caused the locking problem.

3. RFAM adjustment

  • Now I went back to the HWP before the EOM. The HWP was adjusted to minimize the RFAM peak (forgot to record the angle!).
    Also Yaw alignment of the EOM was slightly adjusted to minimize this peak further.
  • The resulting RF leakage of -85dBm was recorded. Shown in the attachment 3. This corresponds to the RIN of 2x10-4.

4. Power supply AC path

  • I noticed that one of the AC tap is obtained the power from the ceiling while another one is taking the power from the wall.
    (Attachment 4)
    Is this what you really want? I am afraid of having big ground loop.
     
  • Suggestion:
    • Get some isolation transformers for the AC / define the grounding point of the experiment.
    • Also ground the optical table and the Al frame.

 

Attachment 1: cabling.jpg  78 kB  Uploaded Thu Nov 18 10:43:04 2010  | Hide | Hide all
cabling.jpg
Attachment 2: reduction.jpg  79 kB  Uploaded Thu Nov 18 10:59:49 2010  | Hide | Hide all
reduction.jpg
Attachment 3: IMG_3714.jpg  85 kB  Uploaded Thu Nov 18 11:23:43 2010  | Hide | Hide all
IMG_3714.jpg
Attachment 4: IMG_3715.jpg  70 kB  Uploaded Thu Nov 18 11:38:25 2010  | Hide | Hide all
IMG_3715.jpg
ELOG V3.1.3-